OpenStudy (anonymous):

By Pmi prove that 1/2+1/4......1/[2^(n)]=1-1/[2^(n)]

OpenStudy (gorv):

induction ???

OpenStudy (anonymous):

I have solved till p(k+1)= 1-1/(2^(k)) +1/(2^[(k+1)]

OpenStudy (anonymous):

Ya induction.

OpenStudy (gorv):

solved for n=1 and n=k ????

OpenStudy (anonymous):

Yes

OpenStudy (gorv):

for n=k what you got ??

OpenStudy (anonymous):

1-1/2^(k)

OpenStudy (gorv):

and for k+1 what will we get ??

OpenStudy (anonymous):

We should prove n=k+1 is 1-1/[2^(k+1)]

OpenStudy (gorv):

first put n=k+1

OpenStudy (gorv):

we have to bring left side equal to right side

OpenStudy (gorv):

put n=k+1 on left

OpenStudy (gorv):

it will form a geomatric progression

OpenStudy (anonymous):

Ya.... The lhs is 1-1/(2^(k)) +1/(2^[(k+1)].... I have to prove that this is equal to1-1/[2^(k+1)].

OpenStudy (gorv):

1/2 +1/4 + 1/8 +1/16+1/32+..................

OpenStudy (gorv):

upto k+1 term

OpenStudy (anonymous):

Ya... That would be 1/2^(k+1)

OpenStudy (gorv):

aum has also a good approch (if dont know abt GP)

OpenStudy (anonymous):

I know GP. And aum's approach is what I generally do. Was confused with solving the exponents.

OpenStudy (aum):

Assume it is true for k. \(\Large p(k) = 1 - \frac{1}{2^k} \) The \((k+1)^{th}\) term is: \(\Large \frac{1}{2^{k+1}}\) Therefore, \(\Large p(k+1) = p(k) + \frac{1}{2^{k+1}}\) \(\Large p(k+1) = 1 - \frac{1}{2^k} + \frac{1}{2^{k+1}} = 1 - \frac{1}{2^k} + \frac{1}{2~*~2^{k}} = \\ \Large 1 - \frac{1}{2^k}(1 - \frac 12) = 1 - \frac{1}{2^k}\frac 12 = 1 - \frac{1}{2^{k+1}} \) If it is true for k, it is true for (k+1). And it is true for n = 1 and therefore it is true for all n.

OpenStudy (anonymous):

Thanks a lot @aum

OpenStudy (aum):

You are welcome.